OpenStudy (anonymous):

10/radical 50 - radical 75/2

OpenStudy (mertsj):

\[\frac{10}{\sqrt{50}}-\frac{\sqrt{75}}{2}=\frac{10}{\sqrt{25(2)}}-\frac{\sqrt{25(3)}}{2}\]

OpenStudy (anonymous):

\[\frac{10}{\sqrt{50}}-\frac{\sqrt{75}}{2}=\sqrt{2}-\frac{5 \sqrt{3}}{2} \]